How much should you pay for a share of stock that offers a constant-growth rate of 10%, requires a 16% rate of return, and is expected to sell for $50 one year from now?
A. $42.00
B. $45.00
C. $45.45
D. $47.00
Answer #1
C. $45.45
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
D2 = (ke-g)P1
D2 = (16%-10%)*50
D2 = (6%)*50
D2 = 3
D1 = 3/1.10                    2.7273
P0 =D1/(ke-g)
P0 =2.7273/(16%-10%)
P0 =2.7273/6%
P0 = 45.45